The height s of a baseball in feet after t seconds is given by s(t)=-16t^2+96t+2
After how many seconds does the baseball reach its maximum height? Show your work.
Find the maximum hieght of the baseball. Show your work.

Answer by Alan3354(69443)

The height s of a baseball in feet after t seconds is given by s(t)=-16t^2+96t+2
After how many seconds does the baseball reach its maximum height?
------------------
The max height is the vertex of the parabola.
Vertex at t = -b/2a = -96/-32 = 3 seconds.
--------------
Max height is s(3)
s(3) = -16*9 + 96*3 + 2
= 146 feet